OpenStudy (watchmath):

Problem from someone here (I forgot who) Compute \[\int_1^\infty \left(\frac{\ln x}{x}\right)^{2001}\]

OpenStudy (watchmath):

Let's make it 2011 \[\int_1^\infty \left(\frac{\ln x}{x}\right)^{2011}\]

OpenStudy (anonymous):

put lnx=y

OpenStudy (cruffo):

if y = ln(x) then e^y = x. change limits when x = 1, y = 0 and when x = infty, y = infty

OpenStudy (cruffo):

so \[\int_0^\infty y^ke^{-ky} dy\]

OpenStudy (anonymous):

how can it be infinity....lnx/x always <1

OpenStudy (cruffo):

looks like a gamma distribution

OpenStudy (watchmath):

Good cruffo :)

OpenStudy (anonymous):

im a noob. wow like a week out from BC Calc and I forget everything :(

OpenStudy (cruffo):

woho!

OpenStudy (watchmath):

knowing gamma function will help alot but if you don't know it you can still solve the problem though.

OpenStudy (cruffo):

yah, by parts - just need to figure out the pattern.

OpenStudy (cruffo):

By parts and looking for a pattern in the outcomes, \[\frac{(k-1)!}{k^k}\]

OpenStudy (anonymous):

yes..gamma functions are like that...

OpenStudy (anonymous):

this is not actually gamma type because e^-ky is here..in gamma the k=1

OpenStudy (cruffo):

I think I remember do all sorts of fancy footwork to change the variables so you could use the distribution and call it a day. Been a few years - have to check my notes!